Lesson eLearning

In a community of 1500 couples, 80 % of couples use male condoms correctly and 20 % incorrectly. The perfect‑use failure rate of condoms is 1.5 % per year and the typical‑use failure rate is 15 % per year.

hard 4 marks Class 12 · Biology · Reproductive Health
Answer
63 pregnancies expected; 41 pregnancies avoided

The correct answer is 63 pregnancies expected; 41 pregnancies avoided — Pregnancies avoided = (63 - 22.5 = 40.5) ≈ 41.

Step-by-step solution

  1. Compute weighted failure probability: - Correct use: (0.8 \times 0.015 = 0.012). - Incorrect use: (0.2 \times 0.15 = 0.03). - Total weighted failure = (0.012 + 0.03 = 0.042).
  2. Multiply by number of couples for expected pregnancies: (1500 \times 0.042 = 63).
  3. If all couples used condoms perfectly, failure = (0.015 \times 1500 = 22.5) ≈ 22.
  4. Pregnancies avoided = (63 - 22.5 = 40.5) ≈ 41.
Common mistake: Using the same failure rate for all couples instead of weighting by correct/incorrect usage

Concept tested: Weighted contraceptive failure rate

Published · How this solution was prepared and checked

← All Reproductive Health questions